OpenStudy (anonymous):

Which equation represents the axis of symmetry of the function y = –2x² + 4x –6

OpenStudy (apoorvk):

for a quadratic ax^2+bx+c=0, vertex is at x=-b/2a, and also, x=-b/2a is the line that is axis of symmetry. calculate and enjoy!

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

i'm not good

OpenStudy (anonymous):

Match up: y = –2x² + 4x – 6 y = ax² + bx + c Hint a = -2, b = .. Can you continue now?